# SOLAR 10.7B: Scaling Large Language Models with Simple yet Effective Depth Up-Scaling

### v2

@@ Abstract
$*{}^{*}$qual Contribution ${}^{\dagger}$ Corresponding Author
@@ 1 Introduction
The field of natural language processing (NLP) has been significantly transformed by the introduction of large language models (LLMs), which have enhanced our understanding and interaction with human language Zhang et al. (2023a). These advancements bring challenges such as the increased need to train ever larger models Rae et al. (2021); Wang et al. (2023); Pan et al. (2023); Lian (2023); Yao et al. (2023); Gesmundo and Maile (2023) owing to the performance scaling law Kaplan et al. (2020); Hernandez et al. (2021); Anil et al. (2023); Kaddour et al. (2023). To efficiently tackle the above, recent works in scaling language models such as a mixture of experts (MoE) Shazeer et al. (2017); Komatsuzaki et al. (2022) have been proposed. While those approaches are able to efficiently and effectively scale-up LLMs, they often require non-trivial changes to the training and inference framework Gale et al. (2023), which hinders widespread applicability. Effectively and efficiently scaling up LLMs whilst also retaining the simplicity for ease of use is an important problem Alberts et al. (2023); Fraiwan and Khasawneh (2023); Sallam et al. (2023); Bahrini et al. (2023).
Inspired by  Komatsuzaki et al. (2022), we present depth up-scaling (DUS), an effective and efficient method to up-scale LLMs whilst also remaining straightforward to use. DUS consists of scaling the base model along the depth dimension and continually pretraining the scaled model. Unlike Komatsuzaki et al. (2022), DUS does not scale the model using MoE and rather use a depthwise scaling method analogous to  Tan and Le (2019) which is adapted for the LLM architecture. Thus, there are no additional modules or dynamism as with MoE, making DUS immediately compatible with easy-to-use LLM frameworks such as HuggingFace Wolf et al. (2019) with no changes to the training or inference framework for maximal efficiency. Furthermore, DUS is applicable to all transformer architectures, opening up new gateways to effectively and efficiently scale-up LLMs in a simple manner. Using DUS, we release SOLAR 10.7B, an LLM with 10.7 billion parameters, that outperforms existing models like Llama 2 Touvron et al. (2023) and Mistral 7B Jiang et al. (2023) in various benchmarks.
@@ 3 Training Details
In the instruction tuning stage, the model is trained to follow instructions in a QA format Zhang et al. (2023b). We mostly use open-source datasets but also synthesize a math QA dataset to enhance the model’s mathematical capabilities. A rundown of how we crafted the dataset is as follows. First, seed math data are collected from the Math Hendrycks et al. (2021) dataset only, to avoid contamination with commonly used benchmark datasets such as GSM8K Cobbe et al. (2021). Then, using a process similar to MetaMath Yu et al. (2023), we rephrase the questions and answers of the seed math data. We use the resulting rephrased question-answer pairs as a QA dataset and call it ‘Synth. Math-Instruct‘.
@@
Table 2: Evaluation results for SOLAR 10.7B and SOLAR 10.7B-Instruct along with other top-performing models. We report the scores for the six tasks mentioned in Sec. 4.1 along with the H6 score (average of six tasks). We also report the size of the models in units of billions of parameters. The type indicates the training stage of the model and is chosen from {Pretrained, Instruction-tuned, Alignment-tuned}. Models based on SOLAR 10.7B are colored purple. The best scores for H6 and the individual tasks are shown in bold.
@@
In the alignment tuning stage, the instruction-tuned model is further fine-tuned to be more aligned with human or strong AI (e.g., GPT4 OpenAI (2023)) preferences using direct preference optimization (DPO) Rafailov et al. (2023). Similar to the instruction tuning stage, we use mostly open-source datasets but also synthesize a math-focused alignment dataset utilizing the ‘Synth. Math-Instruct‘ dataset mentioned in the instruction tuning stage.
@@ 4.1 Experimental Details
We reformatted the instruction datasets with an Alpaca-styled chat template. For datasets such as OpenOrca, which are derived from FLAN Longpre et al. (2023), we filter data that overlaps with the benchmark datasets (see Tab. 8 in Appendix. C for more information). The alignment datasets are in the {prompt, chosen, rejected} triplet format. We preprocess the alignment datasets following Zephyr Tunstall et al. (2023).
@@
In the HuggingFace Open LLM Leaderboard Beeching et al. (2023), six types of evaluation methods are presented: ARC Clark et al. (2018), HellaSWAG Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2022), Winogrande Sakaguchi et al. (2021), and GSM8K Cobbe et al. (2021). We utilize these datasets as benchmarks for evaluation and also report the average scores for the six tasks, e.g., H6.
@@ 4.3 Ablation Studies
We present ablation studies for both the instruction and alignment tuning stages.
@@ 4.3.2 Alignment Tuning
As we utilize DPO for practical alignment tuning, there are additional aspects to ablate such as the SFT base models used. Thus, we present ablations for the different training datasets used for training, the different SFT base models to initialize the DPO model, and finally, the model merging strategy to obtain the final alignment-tuned model.
@@ 5 Conclusion
We introduce SOLAR 10.7B and its fine-tuned variant SOLAR 10.7B-Instruct, which are depth up-scaled (DUS) models with 10.7 billion parameters. They show superior performance over models like Llama 2, Mistral 7B, and Mixtral-7B-Instruct in essential NLP tasks while maintaining computational efficiency. Thus, DUS is effective in scaling-up highly performant LLMs from smaller ones. With more exploration, DUS could be further improved, paving a new path to efficiently scaling LLMs.
@@ Acknowledgements
We would like to extend our gratitude to the teams at Hugging Face, particularly Clémentine Fourrier, Lewis Tunstall, Omar Sanseviero, and Philipp Schmid. Our appreciation also extends to the teams at AWS, notably Ritesh Vajaria, Gal Oshri, Jay Kwon, Brandon Lee, Effie Bae, and Rahul Sharma. We are grateful to the teams at Korea Telecom (KT), especially Jin Hyoung Lee, Jungsuk Park, Sungjoon Park, Hong-rae Wang, Kyeongsoo Jung, and Sunyoong Yoon, whose significant support has been instrumental in ensuring the broad compatibility of our model. Additionally, we would like to extend our thanks to the open community for their invaluable contributions and feedback.
@@ Appendix A Contributions
Dahyun Kim, Chanjun Park, Sanghoon Kim, and Wonsung Lee contributed equally to this paper. Sanghoon Kim led the Foundation Model part, with Dahyun Kim, Wonho Song, Yunsu Kim, and Hyeonwoo Kim. Chanjun Park led the Data and Evaluation (Data-Centric LLM) part, with Yungi Kim, Jihoo Kim, Changbae Ahn, Seonghoon Yang, Sukyung Lee, and Hyunbyung Park. Wonsung Lee led the Adaptation Modeling part, with Gyoungjin Gim, Hyeonju Lee, and Mikyoung Cha. Hwalsuk Lee performed the role of the overall project operation. All these individuals contributed to the creation of SOLAR 10.7B.
@@ B.4 Instruction Tuning
Before instruction tuning, existing methods faced challenges in effectively guiding and controlling the behavior of large language models Zhang et al. (2023b). The sheer complexity of these models made it difficult to ensure precise and task-oriented responses. The need for a more targeted approach arose from the limitations of existing methods, leading to the development of instruction tuning. This targeted approach enables better control over the model’s behavior, making it more suitable for specific tasks and improving its overall performance in alignment with user-defined objectives. Therefore, instruction tuning is computationally efficient and facilitates the rapid adaptation of LLMs to a specific domain without requiring extensive retraining or architectural changes.

### v3

@@ Abstract
Equal Contribution Corresponding Author
@@ 1 Introduction
The field of natural language processing (NLP) has been significantly transformed by the introduction of large language models (LLMs), which have enhanced our understanding and interaction with human language Zhao et al. (2023). These advancements bring challenges such as the increased need to train ever larger models Rae et al. (2021); Wang et al. (2023); Pan et al. (2023); Lian (2023); Yao et al. (2023); Gesmundo and Maile (2023) owing to the performance scaling law Kaplan et al. (2020); Hernandez et al. (2021); Anil et al. (2023); Kaddour et al. (2023). To efficiently tackle the above, recent works in scaling language models such as a mixture of experts (MoE) Shazeer et al. (2017); Komatsuzaki et al. (2022) have been proposed. While those approaches are able to efficiently and effectively scale-up LLMs, they often require non-trivial changes to the training and inference framework Gale et al. (2023), which hinders widespread applicability. Effectively and efficiently scaling up LLMs whilst also retaining the simplicity for ease of use is an important problem Alberts et al. (2023); Fraiwan and Khasawneh (2023); Sallam et al. (2023); Bahrini et al. (2023).
Inspired by  Komatsuzaki et al. (2022), we present depth up-scaling (DUS), an effective and efficient method to up-scale LLMs whilst also remaining straightforward to use. DUS consists of scaling the number of layers in the base model and continually pretraining the scaled model. Unlike Komatsuzaki et al. (2022), DUS does not scale the model using MoE and rather use a depthwise scaling method analogous to  Tan and Le (2019) which is adapted for the LLM architecture. Thus, there are no additional modules or dynamism as with MoE, making DUS immediately compatible with easy-to-use LLM frameworks such as HuggingFace Wolf et al. (2019) with no changes to the training or inference framework for maximal efficiency. Furthermore, DUS is applicable to all transformer architectures, opening up new gateways to effectively and efficiently scale-up LLMs in a simple manner. Using DUS, we release SOLAR 10.7B, an LLM with 10.7 billion parameters, that outperforms existing models like Llama 2 Touvron et al. (2023) and Mistral 7B Jiang et al. (2023) in various benchmarks.
@@ 3 Training Details
In the instruction tuning stage, the model is trained to follow instructions in a QA format Zhang et al. (2023). We mostly use open-source datasets but also synthesize a math QA dataset to enhance the model’s mathematical capabilities. A rundown of how we crafted the dataset is as follows. First, seed math data are collected from the Math Hendrycks et al. (2021) dataset only, to avoid contamination with commonly used benchmark datasets such as GSM8K Cobbe et al. (2021). Then, using a process similar to MetaMath Yu et al. (2023), we rephrase the questions and answers of the seed math data. We use the resulting rephrased question-answer pairs as a QA dataset and call it ‘Synth. Math-Instruct‘.
@@
Table 2: Evaluation results in the Open LLM Leaderboard for SOLAR 10.7B and SOLAR 10.7B-Instruct along with other top-performing models. We report the scores for the six tasks mentioned in Sec. 4.1 along with the H6 score (average of six tasks). We also report the size of the models in units of billions of parameters. The type indicates the training stage of the model and is chosen from {Pretrained, Instruction-tuned, Alignment-tuned}. Models based on SOLAR 10.7B are colored purple. The best scores for H6 and the individual tasks are shown in bold.
@@
In the alignment tuning stage, the instruction-tuned model is further fine-tuned to be more aligned with human or strong AI (e.g., GPT4 OpenAI (2023)) preferences using sDPO Kim et al. (2024a), an improved version of direct preference optimization (DPO) Rafailov et al. (2023). Similar to the instruction tuning stage, we use mostly open-source datasets but also synthesize a math-focused alignment dataset utilizing the ‘Synth. Math-Instruct‘ dataset mentioned in the instruction tuning stage.
@@ 4.1 Experimental Details
We reformatted the instruction datasets with an Alpaca-styled chat template. For datasets such as OpenOrca, which are derived from FLAN Longpre et al. (2023), we filter data that overlaps with the benchmark datasets (see Tab. 8 in Appendix. C for more information). The alignment datasets are in the {prompt, chosen, rejected} triplet format. We preprocess the alignment datasets following Zephyr Tunstall et al. (2023). We use Dataverse Park et al. (2024) for data preprocessing.
@@
In the HuggingFace Open LLM Leaderboard Beeching et al. (2023), six types of evaluation methods are presented: ARC Clark et al. (2018), HellaSWAG Zellers et al. (2019), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2022), Winogrande Sakaguchi et al. (2021), and GSM8K Cobbe et al. (2021). We utilize these datasets as benchmarks for evaluation and also report the average scores for the six tasks, e.g., H6. We either submit directly to the Open LLM Leaderboard or utilize Evalverse Kim et al. (2024b) for running evaluations locally.
@@ 4.3 Ablation Studies
We present ablation studies for both the instruction and alignment tuning stages. Note that the evaluation results for the following studies are ran locally and may vary from results obtained by submitting to the Open LLM Leaderboard.
@@ 4.3.2 Alignment Tuning
As we utilize sDPO for practical alignment tuning, there are additional aspects to ablate such as the SFT base models used. Thus, we present ablations for the different training datasets used for training, the different SFT base models to initialize the sDPO training, and finally, the model merging strategy to obtain the final alignment-tuned model.
@@ 5 Conclusion
We introduce SOLAR 10.7B and its fine-tuned variant SOLAR 10.7B-Instruct, which are depth up-scaled (DUS) models with 10.7 billion parameters [^4] Preprint version is available on https://arxiv.org/abs/2312.15166.. They show superior performance over models like Llama 2, Mistral 7B, and Mixtral-7B-Instruct in essential NLP tasks while maintaining computational efficiency. Thus, DUS is effective in scaling-up highly performant LLMs from smaller ones. With more exploration, DUS could be further improved, paving a new path to efficiently scaling LLMs.
@@ Acknowledgements
We would like to extend our gratitude to the teams at Hugging Face, particularly Clémentine Fourrier, Lewis Tunstall, Omar Sanseviero, and Philipp Schmid. Our appreciation also extends to the teams at AWS, notably Rahul Sharma, Jeongwon Yoon, Nieves Garcia, Ritesh Vajaria, Gal Oshri, Jay Kwon, Brandon Lee and Effie Bae. We are grateful to the teams at Korea Telecom (KT), especially Jin Hyoung Lee, Jungsuk Park, Sungjoon Park, Hong-rae Wang, Kyeongsoo Jung, and Sunyoong Yoon, whose significant support has been instrumental in ensuring the broad compatibility of our model. Additionally, we would like to extend our thanks to the open community for their invaluable contributions and feedback.
@@ Appendix A Contributions
Sanghoon Kim, Dahyun Kim, Chanjun Park, Wonsung Lee, Wonho Song, Yunsu Kim and Hyeonwoo Kim contributed equally to this paper. Sanghoon Kim led the Foundation Model part, with Dahyun Kim, Wonho Song, Yunsu Kim, and Hyeonwoo Kim. Chanjun Park led the Data and Evaluation (Data-Centric LLM) part, with Yungi Kim, Jihoo Kim, Changbae Ahn, Seonghoon Yang, Sukyung Lee, and Hyunbyung Park. Wonsung Lee led the Adaptation Modeling part, with Gyoungjin Gim, Hyeonju Lee, and Mikyoung Cha. Hwalsuk Lee performed the role of the overall project operation. Dahyun Kim and Chanjun Park were the main technical writers. All these individuals contributed to the creation of SOLAR 10.7B.
@@ B.4 Instruction Tuning
Before instruction tuning, existing methods faced challenges in effectively guiding and controlling the behavior of large language models Zhang et al. (2023). The sheer complexity of these models made it difficult to ensure precise and task-oriented responses. The need for a more targeted approach arose from the limitations of existing methods, leading to the development of instruction tuning. This targeted approach enables better control over the model’s behavior, making it more suitable for specific tasks and improving its overall performance in alignment with user-defined objectives. Therefore, instruction tuning is computationally efficient and facilitates the rapid adaptation of LLMs to a specific domain without requiring extensive retraining or architectural changes.